Find the total heat produced in the loop of the previous problem during the interval 0 to 30 s if the resistance of the loop is 4.5 mΩ.

Grade:11

1 Answers

Jitender Pal
9 years ago
Sol. As heat produced is a scalar prop. So, net heat produced = H base a + H base b + H base c + H base d R = 4.5 mΩ = 4.5 ×10–3 Ω a) e = 3 × 10^–4 V i = e/R = 3x 10^-4 / 4.5 x 10^-3 = 6.7 x 10^-2 Amp. H base a = (6.7 ×10^–2)^2 × 4.5 × 10^–3 × 5 H base b = H base d = 0 [since emf is induced for 5 sec] H base c = (6.7 ×10^–2)^2 × 4.5 × 10^–3 × 5 So Total heat = H base a + H base c = 2 × (6.7 ×10^–2)^2 × 4.5 × 10^–3 × 5 = 2 × 10^–4 J.